find the equation of the line passing through point (4,-1) and parallel to the line joining A(-2,5) and B(4,9)
Answer by oscargut(2103)

slope of the line AB where A(-2,5) and B(4,9) is (9-5)/(4+2) = 4/6 = 2/3

y = (2/3)x+b
Since (4,-1) is in the line:
-1 = (2/3)4+b
b = -1 -8/3
b = -11/3
Answer: y = (2/3)x -11/3
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com